OpenStudy (anonymous):

Medal and fan 4. Directions: Type the correct answer in each box. Spell all words correctly, and use numerals instead of words for numbers. If necessary, use decimals instead of fractions. Two brothers, Ty and Cam, went to a local bookstore to buy books for school. Ty bought 8 books for $113.92, and Cam bought 6 books for $92.82. They then compared their average cost per book.

OpenStudy (anonymous):

Try taking a screenshot and posting it

OpenStudy (anonymous):

Ty paid an average cost of $ (blank)per book. Cam paid an average cost of $ ( blank) per book. So, paid an average cost of $ (blank) more per book than his brother.

OpenStudy (michaelac959):

add it all together and divide by the number of books.

OpenStudy (anonymous):

Wait let me start over, it was wrong

OpenStudy (anonymous):

@Superwholock4445 did you get my message?

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

Ty paid an average cost of $(blank) per book. Cam paid an average cost of $(blank) per book. So, (blank) paid an average cost of $ ( blank) more per book than his brother.

OpenStudy (michaelac959):

so 113.92+92.82=206.74 then 206.74/14=14.76

OpenStudy (anonymous):

Ill come back to this, I have some questions of my own to ask

OpenStudy (michaelac959):

sorry i did it wrong let me start over.

OpenStudy (anonymous):

K

OpenStudy (michaelac959):

8/113.92=14.24 92.82/6=15.47

OpenStudy (michaelac959):

Ty paid an average cost of $14.24 per book. Cam paid an average cost of $15.47 per book. So, cam paid an average cost of $1.23more per book than his brother.

OpenStudy (anonymous):

Can you meet me on my next page please?